### v1

@@ Abstract
To realize effective large-scale, real-world robotic applications, we must evaluate how well our robot policies adapt to changes in environmental conditions. Unfortunately, a majority of studies evaluate robot performance in environments closely resembling or even identical to the training setup. We present The Colosseum, a novel simulation benchmark, with 20 diverse manipulation tasks, that enables systematical evaluation of models across 12 axes of environmental perturbations. These perturbations include changes in color, texture, and size of objects, table-tops, and backgrounds; we also vary lighting, distractors, and camera pose. Using The Colosseum, we compare 4 state-of-the-art manipulation models to reveal that their success rate degrades between 30-50% across these perturbation factors. When multiple perturbations are applied in unison, the success rate degrades $\geq$75%. We identify that changing the number of distractor objects, target object color, or lighting conditions are the perturbations that reduce model performance the most. To verify the ecological validity of our results, we show that our results in simulation are correlated ($\bar{R}^{2}=0.614$) to similar perturbations in real-world experiments. We open source code for others to use The Colosseum, and also release code to 3D print the objects used to replicate the real-world perturbations. Ultimately, we hope that The Colosseum will serve as a benchmark to identify modeling decisions that systematically improve generalization for manipulation.
@@ I Introduction
Figure 1: Evaluating generalization with The Colosseum. Task-averaged success rate for 4 SotA behavior cloning models over 12 perturbation factors and 20 robotic manipulation tasks. Changes in RGB input space affects all models due to end-to-end RGB-based training. Image-based models are also affected by camera pose change, while models without in-the-wild pretraining suffer in the presence of distractors.
Figure 2: The Colosseum Challenge. This challenge is designed to enhance generalization of Behavior Cloning (BC) models in robotic manipulation tasks. It involves four key phases: 1) Participants generate a standard training dataset from 20 tasks with 100 demonstrations each, without perturbation_factors. 2) Participants train their BC models using this standardized dataset. 3) The models are restricted to evaluate over a fixed 25 episodes across 12 different perturbation_factors. 4) Models are ranked on a leaderboard based on the percentage change in their performance across these factors. We’ve shown that simulation aligns with real-world evaluation, so participants can expect similar generalization when participating in the simulation benchmark.
@@
We introduce The Colosseum, a comprehensive benchmark aimed at systematically evaluating the generalization of robot manipulation to environmental perturbations. The Colosseum introduces perturbations across 20 different tasks from the RLBench [29] framework, spanning 12 dimensions of perturbations. These perturbations include object color, object texture, object size, table color, and table texture, the presence of distractor objects, and changes to the camera pose. The Colosseum also includes a parallel real world evaluation with task setups and objects reproducible via open-sourced 3D printing models.
@@ II-A Robotic Manipulation Benchmarks
Benchmarks in computer vision have significantly advanced the development of more generalized vision systems in recent decades by introducing numerous challenges and leaderboards [14, 32, 34], subsequently scaling into extensive foundational vision models [7, 3, 35]. Similarly, robotics datasets have demonstrated considerable diversity across various dimensions [12, 17, 31], particularly with the evolution of imitation learning and, more specifically, behavior cloning (BC). This progress has led to a proliferation of datasets and benchmarks aimed at assessing BC model task performance. Furthermore, most of these robotic benchmarks focus on evaluating model’s capability to adapt to new tasks by changing the nature of the task [29, 68], it’s functionalities[23, 36], or even environment [13, 48]. However, a gap remains in systematically evaluating and comparing different BC models on a large-scale, both in simulation and in real-world reproducibly.
Factor World [62] and KitchenShift [63] are similar efforts to The Colosseum, but have 11 and 7 perturbation types, by contrast to the 12 over 20 tasks in The Colosseum. Additionally, The Colosseum supports both 2D and 3D models, unlike Factor World which only evaluates on 2D visual-motor policies. Furthermore, we employed 3D printed objects to test with a Franka Panda robot arm to enable easy replication of our real-world experiments.
@@ II-B Robotic Manipulation Methods
Figure 3: The Colosseum benchmark distribution. This benchmark encompasses 12 perturbation_factors within 20 distinct RLBench tasks, categorized into three tiers (simple, intermediate, and complex) according to the number of way-points involved (task horizon). Collectively, The Colosseum presents 17,431 unique task perturbations instances.
@@ II-C Generalization in Robotic Manipulation
Traditionally, enhancing generalization in imitation learning involves employing image or 3D data augmentation techniques, akin to those used in computer vision. These techniques encompass random shifts, color adjustments, and rotations [57, 33, 22, 64]. Sim-to-real transfer methods utilize advanced simulation environments for domain randomization and preliminary policy training before real-world application [43, 24]. Recent developments in extensive vision and language models present a novel path to generalization: pretrained on vast image and language datasets, these models offer potentially more robust representations for robotic manipulation or can even directly inform actions [1, 25, 9, 5].
@@ III The Colosseum
The Colosseum is a comprehensive simulation benchmark, built by extending RLBench, consisting of 20 diverse robotic manipulation tasks, each enabled with 12 perturbation_factors. We base off of RLBench as it provides a variety of realistically useful tasks, with a scripted demonstration generation framework, broad variance in their task horizon (i.e. the number of controller steps required to complete the task) and primitive actions (such as pick, place, open, close, turn, and slide). We define perturbation_factors as scene properties, such as object color or lighting conditions. These properties can be changed to cause data distribution shifts at test-time such that the input distribution changes $p(x_{test})\neq p(x_{train})$, but the conditional probability of action distribution remains the same $p(y_{test}|x_{test})=p(y_{train}|x_{train})$, as the underlying task does not change. This form of distribution shift in Out-of-Distribution (OoD) generalization research is referred to as covariate shift [26].
With The Colosseum, we proposed over 17,431 unique task instances from a list of 20 tasks. The tasks are also categorized into three tiers of difficulties based on the task horizon, which inherently makes the tasks harder due to compounding error in BC [52]. The detailed breakdown of the number of unique instance per variation is also shown in Figure 3.
@@ III-B Perturbation Factors
We create 12 perturbation_factors and apply each of them to the above 20 tasks where compatible. We categorize them as follows:
@@ III-C Implementation of Perturbation Factors
Following RLBench, our implementation utilizes PyRep [28], a low-level API, to interact with the underlying CoppeliaSim [51] simulator. The PyRep API allows the control of simulator properties such as color, texture, scaling, and pose of the objects in the scene. We implement 12 perturbation_factors (shown in Figure 2: part 3) as an extension of the RLBench task benchmark. We expose the configuration of supported perturbations via configuration files written in YAML. Our implementation is easily extensible for other researchers to build on and edit the benchmark, for adding new perturbation_factors or new tasks, along with the ease of configuring any combination of perturbation_factors and their parameters, where compatible.
To implement texture or color perturbations, we randomly sample a texture or color from our curated set. We provide a set of 213 textures and 20 colors. These assets were used for implementing MO_Color, MO_Texture, RO_Color, RO_Texture, Table_Color, Table_Texture, and Background_Texture. To implement size perturbations (MO_Size and RO_Size), we sample a scaling factor from a continuous range, specified for each object that supports this factor. For instance, the range for MO_Size in the task basketball_in_hoop is $[0.75,1.25]$, differs from that in task hockey $[0.95,1.05]$, as this parameter is quite dependent on the conditions of the objects in the scene. We including remaining task’s parameters in the Appendix. The waypoints get re-scaled with the object, and if not, we reposition them proportionally with respect to the object center, while ensuring that RLBench’s scripted demonstration generation from waypoints remains functional. The Distractor objects are sampled from a set of 78 object models taken from YCB Object Dataset [6] and converted to CoppeliaSim compatible .ttm format. We utilize predefined object spawn boundaries to place these objects on table-top or on another object. To modify Light Color, we randomly sample RGB values from our specified range of $[0.0,0.0,0.0]$ to $[0.5,0.5,0.5]$, and apply it to all 3 directional lights surrounding the scene. We perturb Camera Pose for 3 cameras — front, left shoulder, and right shoulder — by changing their positions and orientations in Euler angles, sampled from ranges $[-0.1,-0.1,-0.1]$ to $[0.1,0.1,0.1]$ and $[-0.05,-0.05,-0.05]$ to $[0.05,0.05,0.05]$ respectively. We provide our texture, color, and object model assets with the benchmark code, which can also be augmented easily for additional assets, as required.
@@ III-E The Colosseum Challenge
We propose The Colosseum Challenge to enable development of generalizable Behavior Cloning (BC) models for robotic manipulation tasks. As shown in Figure 2, the challenge involves four key phases: 1) Participants generate a standard training dataset for 20 tasks with 100 demonstrations each, without perturbation_factors. 2) Participants train their BC models using this standardized dataset. 3) The models should evaluate over a fixed 25 episodes set of each of the 12 perturbation_factors. 4) Models are ranked on a leaderboard based on the percentage change in their performance across these perturbation_factors. Demonstrations data for training and testing can be generated via scripted experts from RLBench. Each task is equipped to instantiate object pose-variated episodes ensuring an inexhaustible supply of task-specific demonstrations. Demonstrations are collected automatically through motion planners that navigate through manually defined waypoints.
@@ IV-B Baselines
We study 4 SotA baselines, including two 2D learning models (R3M-MLP, MVP-MLP) and two 3D learning models (PerAct, RVT). We choose these methods as baselines as they establish themselves as strong robot learning methods. They are also diverse in their approach, allowing us to study the effect of aspects such as, pretraining and 2D vs 3D based learning. For all baselines, the language is encoded using a frozen CLIP [49] model.
@@
[…]
@@
IV-C Training and Evaluation Protocol
@@
Figure 5: Task-averaged success rate % change for 4 baseline models on perturbation_factors, compared to No Perturbation test set. We report the evaluation with All Perturbations enabled, followed by each individual factor, average of all individual factors, and on RLBench variations (that is sampled from the same distribution as the training set). The images on top show failure examples for each factor with captions explaining the failure. normal-$\bullet$ indicates undefined value when the corresponding No Perturbation task averages are also 0. $\circ$ indicates 0% change with respect to No Perturbation task average.
IV-D Real-World Setup
@@ V-A Performance of different baselines on The Colosseum
For 3D-models (RVT and PerAct), we observe that the most affecting factors are color-related including object, table and light colors as well as presence of Distractors, while other factors cause a smaller performance decrease. Since RVT and PerAct are both trained end-to-end with RGB images or voxel grid with RGB channels, the color perturbations remain challenging for these models as well. These models lack any real-world pretraining, thus, the presence of Distractors puts the scene out of distribution, significantly affecting their performance. We observe that these model are robust to changes in Camera_Pose, because they do not directly learn on captured view. They instead preprocess the input RGBD views into a voxel grid or re-rendered novel views. For these models, while each factor doesn’t lead to a very significant effect on their performance, all factors combined in one scene (All Perturbations) cause a significant decrease.

### v2

@@ Abstract
To realize effective large-scale, real-world robotic applications, we must evaluate how well our robot policies adapt to changes in environmental conditions. Unfortunately, a majority of studies evaluate robot performance in environments closely resembling or even identical to the training setup. We present The Colosseum, a novel simulation benchmark, with 20 diverse manipulation tasks, that enables systematical evaluation of models across 14 axes of environmental perturbations. These perturbations include changes in color, texture, and size of objects, table-tops, and backgrounds; we also vary lighting, distractors, physical properties perturbations and camera pose. Using The Colosseum, we compare 5 state-of-the-art manipulation models to reveal that their success rate degrades between 30-50% across these perturbation factors. When multiple perturbations are applied in unison, the success rate degrades $\geq$75%. We identify that changing the number of distractor objects, target object color, or lighting conditions are the perturbations that reduce model performance the most. To verify the ecological validity of our results, we show that our results in simulation are correlated ($\bar{R}^{2}=0.614$) to similar perturbations in real-world experiments. We open source code for others to use The Colosseum, and also release code to 3D print the objects used to replicate the real-world perturbations. Ultimately, we hope that The Colosseum will serve as a benchmark to identify modeling decisions that systematically improve generalization for manipulation.
@@ I Introduction
Figure 1: Evaluating generalization with The Colosseum. Task-averaged success rate for 5 SotA robotic manipulation policies over 14 perturbation factors and 20 robotic manipulation tasks. Changes in RGB input space affects all models due to end-to-end RGB-based training. Image-based models are also affected by camera pose change, while models without in-the-wild pretraining suffer in the presence of distractors.
Figure 2: The Colosseum Challenge. This challenge is designed to enhance generalization of Behavior Cloning (BC) models in robotic manipulation tasks. It involves four key phases: 1) Participants generate a standard training dataset from 20 tasks with 100 demonstrations each, without perturbation_factors. 2) Participants train their BC models using this standardized dataset. 3) The models are restricted to evaluate over a fixed 25 episodes across 14 different perturbation_factors. 4) Models are ranked on a leaderboard based on the percentage change in their performance across these factors. We’ve shown that simulation aligns with real-world evaluation, so participants can expect similar generalization when participating in the simulation benchmark.
@@
We introduce The Colosseum, a comprehensive benchmark aimed at systematically evaluating the generalization of robot manipulation to environmental perturbations. The Colosseum introduces perturbations across 20 different tasks from the RLBench [32] framework, spanning 14 dimensions of perturbations. These perturbations include object color, object texture, object size, table color, table texture, the presence of distractor objects, changes to the camera pose, and changes to physical properties like friction and mass. The Colosseum also includes a parallel real world evaluation with task setups and objects reproducible via open-sourced 3D printing models.
@@
Benchmark Simulator No. of perturbations No. of tasks Physical perturbation Real-world reproducibility No. of SoTA models
GROOT [78] LIBERO [39] 3 3 $\times$ $\checkmark$ 4
VLMBench [77] RLBench [32] 8 8 $\times$ $\times$ 3
KitchenShift [71] Isaac Sim [42] 7 3 $\times$ $\times$ 5
FactorWorld [70] MuJoCo [68] 11 19 $\times$ $\times$ 2
The Colosseum (ours) RLBench [32] $14^{*}$ 20 $\checkmark$ $\checkmark$ 5
TABLE I: Generalization benchmarks comparison. The Colosseum is the largest and most diverse benchmark for evaluating generalization in robotic manipulation policies, covering a wide range of variations and tasks. It is also the first to incorporate physical property perturbations. Similar to GROOT [78], we offer comprehensive instructions and 3D printed components for replicating real-world experiments. Additionally, we evaluated a variety of robot manipulation policies. One asterisk (*) indicates that, unlike the other benchmarks, the object’s position is considered a default perturbation and was not counted as an additional perturbation.
@@ II-A Robotic Manipulation Benchmarks
Benchmarks in computer vision have significantly advanced the development of more generalized vision systems in recent decades by introducing numerous challenges and leaderboards [15, 36, 38], subsequently scaling into extensive foundational vision models [7, 3, 40]. Similarly, robotics datasets have demonstrated considerable diversity across various dimensions [13, 17, 19, 34, 73], particularly with the evolution of imitation learning and, more specifically, behavior cloning (BC). This progress has led to a proliferation of datasets and benchmarks aimed at assessing BC model task performance. Furthermore, most of these robotic benchmarks focus on evaluating model’s capability to adapt to new tasks by changing the nature of the task [32, 79], it’s functionalities[26, 41], or even environment [14, 55]. However, a gap remains in systematically evaluating and comparing different BC models on a large-scale, both in simulation and in real-world. Additionally, many of the benchmarks and datasets with perturbations have been specifically defined or curated as part of various BC or reinforcement learning works [78, 25, 73]. However, they do not provide an unified framework to evaluate all potential perturbations for generalization, primarily because they are not the main focus of these works.
Factor World [70] and KitchenShift [71] are similar efforts to The Colosseum. However, Factor World encompasses only 11 variation factors across 19 tasks, whereas KitchenShift contains 7 variation factors across 3 tasks. In contrast, The COLOSSEUM boasts 14 factors of variation over 20 tasks. Beyond that, The Colosseum supports both 2D and 3D models, unlike Factor World which only evaluates on 2D visual-motor policies. Furthermore, we employed 3D printed objects to test with a Franka Panda robot arm to enable easy replication of our real-world experiments.
@@ II-B Robotic Manipulation Methods
Figure 3: The Colosseum benchmark distribution. This benchmark encompasses 14 perturbation_factors within 20 distinct RLBench tasks, categorized into three tiers (simple, intermediate, and complex) according to the number of way-points involved (task horizon). Collectively, The Colosseum presents 20,371 unique task perturbations instances.
@@ II-C Generalization in Robotic Manipulation
Traditionally, enhancing generalization in imitation learning involves employing image or 3D data augmentation techniques, akin to those used in computer vision. These techniques encompass random shifts, color adjustments, and rotations [64, 37, 24, 72]. Sim-to-real transfer methods utilize advanced simulation environments for domain randomization and preliminary policy training before real-world application [50, 27, 45]. Recent developments in extensive vision and language models present a novel path to generalization: pretrained on vast image and language datasets, these models offer potentially more robust representations for robotic manipulation or can even directly inform actions [1, 28, 9, 5].
@@ III The Colosseum
The Colosseum is a comprehensive simulation benchmark, built by extending RLBench, consisting of 20 diverse robotic manipulation tasks, each enabled with 14 perturbation_factorsẆe base off of RLBench as it provides a variety of realistically useful tasks, with a scripted demonstration generation framework, broad variance in their task horizon (i.e. the number of controller steps required to complete the task) and primitive actions (such as pick, place, open, close, turn, and slide). We define perturbation_factors as scene properties, such as object color or lighting conditions. These properties can be changed to cause data distribution shifts at test-time such that the input distribution changes $p(x_{test})\neq p(x_{train})$, but the conditional probability of action distribution remains the same $p(y_{test}|x_{test})=p(y_{train}|x_{train})$, as the underlying task does not change. This form of distribution shift in Out-of-Distribution (OoD) generalization research is referred to as covariate shift [29].
With The Colosseum, we proposed over 20,371 unique task instances from a list of 20 tasks. The tasks are also categorized into three tiers of difficulties based on the task horizon, which inherently makes the tasks harder due to compounding error in BC [59]. The detailed breakdown of the number of unique instance per variation is also shown in Figure 3.
@@ III-B Perturbation Factors
We create 14 perturbation_factors and apply each of them to the above 20 tasks where compatible. We categorize them as follows:
@@
Physical perturbation
Factors that affect physical properties of the objects involved in the task, such as, Object_Friction where the task involves sliding of an object, and Object_Mass where the gripper needs to adapt to force required for moving the object.
@@ III-C Implementation of Perturbation Factors
Following RLBench, our implementation utilizes PyRep [31], a low-level API, to interact with the underlying CoppeliaSim [58] simulator. The PyRep API allows the control of simulator properties such as color, texture, scaling, and pose of the objects in the scene. We implement 14 perturbation_factors (shown in Figure 2: part 3) as an extension of the RLBench task benchmark. We expose the configuration of supported perturbations via configuration files written in YAML. Our implementation is easily extensible for other researchers to build on and edit the benchmark, for adding new perturbation_factors or new tasks, along with the ease of configuring any combination of perturbation_factors and their parameters, where compatible.
To implement texture or color perturbations, we randomly sample a texture or color from our curated set. We provide a set of 213 textures and 20 colors. These assets were used for implementing MO_Color, MO_Texture, RO_Color, RO_Texture, Table_Color, Table_Texture, and Background_Texture. To implement size perturbations (MO_Size and RO_Size), we sample a scaling factor from a continuous range, specified for each object that supports this factor. For instance, the range for MO_Size in the task basketball_in_hoop is $[0.75,1.25]$, differs from that in task hockey $[0.95,1.05]$, as this parameter is quite dependent on the conditions of the objects in the scene. We include the remaining task parameters in the Appendix. The waypoints get re-scaled with the object, and if not, we reposition them proportionally with respect to the object center, while ensuring that RLBench’s scripted demonstration generation from waypoints remains functional. The Distractor objects are sampled from a set of 78 object models taken from the YCB Object Dataset [6] and converted to CoppeliaSim compatible .ttm format. We utilize predefined object spawn boundaries to place these objects on table-top or on another object. To modify Light Color, we randomly sample RGB values from our specified range of $[0.0,0.0,0.0]$ to $[0.5,0.5,0.5]$, and apply it to all 3 directional lights surrounding the scene. We perturb Camera Pose for 3 cameras — front, left shoulder, and right shoulder — by changing their positions and orientations in Euler angles, sampled from ranges $[-0.1,-0.1,-0.1]$ to $[0.1,0.1,0.1]$ and $[-0.05,-0.05,-0.05]$ to $[0.05,0.05,0.05]$ respectively. Object_Friction is implemented by changing the friction coefficient of the object with a value sampled from the range $[0.75,1.0]$. Object_Mass changes the mass of objects with a value sampled from a given range, where the ranges are task dependent (provided in Appendix). We provide our texture, color, and object model assets with the benchmark code, which can also be augmented easily for additional assets, as required.
@@ III-E The Colosseum Challenge
We propose The Colosseum Challenge to enable development of generalizable Behavior Cloning (BC) models for robotic manipulation tasks. As shown in Figure 2, the challenge involves four key phases: 1) Participants generate a standard training dataset for 20 tasks with 100 demonstrations each, without perturbation_factors. 2) Participants train their BC models using this standardized dataset. 3) The models should evaluate over a fixed 25 episodes set of each of the 14 perturbation_factors. 4) Models are ranked on a leaderboard based on the percentage change in their performance across these perturbation_factors. Demonstrations data for training and testing can be generated via scripted experts from RLBench. Each task is equipped to instantiate object pose-variated episodes ensuring an inexhaustible supply of task-specific demonstrations. Demonstrations are collected automatically through motion planners that navigate through manually defined waypoints.
@@ IV-B Baselines
We study 5 SotA baselines, including one zero-shot open vocabulary model (VoxPoser), two 2D learning models (R3M-MLP, MVP-MLP) and two 3D learning models (PerAct, RVT). We choose these methods as baselines as they establish themselves as strong robot learning methods. They are also diverse in their approach, allowing us to study the effect of aspects such as, pretraining and 2D vs 3D based learning. For all baselines, the language is encoded using a frozen CLIP [56] model.
@@
PerAct is a transformer-based robotic manipulation BC model that takes tokenized voxel grid and language instruction as the input, to predict discretized voxel grid translation point, discrete rotation in Euler angles, and gripper’s binary (open/close) state. PerAct works with 3D voxel grid tokens, akin to visual patch tokens or language tokens in vision or language transformers. Following the original implementation, we use a voxel grid of size $100^{3}$, corresponding to an actual volume of 1.0m3. The patch tokens of size $5^{3}$ are encoded via a 3D convolution layer with kernel-size and stride of 5, resulting in $800020^{3}=$ 8000 voxel observation tokens. Actions are discretized via voxelized keypoint-based action prediction. The actions are then predicted as the next-best voxel that is closest to the center of the gripper fingers for the next translation pose. Rotation pose is discretized into bins of $5^{o}$ increments. The input to the model is the encoded language instruction, proprioception, and 4 camera RGBD views, which gets preprocessed into a voxel grid with voxel occupancy and RGB channels. The model has $\sim$33M trainable parameters. We train this model with batch size 16 for 300k training iterations.
@@
IV-C Zero-Shot manipulation model using Large Pretrained World Models
VoxPoser[28] is a formulation that aims to extract affordances and constraints using LLMs. Through code, it composes 3D value maps in observation space to guide robotic interactions. We utilized their RLBench implementation for VoxPoser, providing variation descriptions from each demonstration as input language text. We manually annotated all corresponding RLBench[32] objects with their respective object names mentioned in the variation descriptions. We conducted a zero-shot evaluation of VoxPoser using The Colosseum’s evaluation protocol, without any training involved.
IV-D Training and Evaluation Protocol
@@
Figure 5: Task-averaged success rate % change for 4 baseline models on perturbation_factors, compared to No Perturbation test set. We report the evaluation with All Perturbations enabled, followed by each individual factor, average of all individual factors, and on RLBench variations (that is sampled from the same distribution as the training set). The images on top show failure examples for each factor with captions explaining the failure. $\bullet$ indicates undefined value when the corresponding No Perturbation task averages are also 0. $\circ$ indicates 0% change with respect to No Perturbation task average.
IV-E Real-World Setup
@@ V-A Performance of different baselines on The Colosseum
For zero-shot manipulation models using Large Pretrained World Models, we observed that the system demonstrates robust generalization capabilities across various conditions, particularly excelling in tasks where it is predisposed to succeed. Specifically, for the two tasks in which VoxPoser excels, it maintains consistent performance across all variants. For example, in the task slide_block_to_target, the performance difference between the No Perturbation scenario and the average of all perturbations is a mere 3.21% relative to the No Perturbation performance. This aligns with our expectation that leveraging large pretrained world models enables the recognition of significant changes in environments and object perturbations.
For 3D-models (RVT and PerAct), we observe that the most affecting factors are color-related including object, table and light colors as well as presence of Distractors, while other factors cause a smaller performance decrease. Since RVT and PerAct are both trained end-to-end with RGB images or voxel grid with RGB channels, the color perturbations remain challenging for these models as well. These models lack any real-world pretraining, thus, the presence of Distractors puts the scene out of distribution, significantly affecting their performance. We observe that these model are robust to changes in Camera_Pose, because they do not directly learn on captured view. They instead preprocess the input RGBD views into a voxel grid or re-rendered novel views. For these models, while each factor doesn’t lead to a very significant effect on their performance, all factors combined in one scene (All Perturbations) cause a significant decrease. On physical perturbations RVT performs better than PerAct, perhaps because modelling in RVT is more robust for keypoint prediction than PerAct under these perturbations. Physical perturbation results for other models are inconlcusive as they cannot perform the tasks that support these perturbations.